OpenStudy (anonymous):

2. The following formula relates three quantities: Force (F), mass (m), and acceleration (a). F = ma a. Solve this equation for a. b. If the force is F = –25 units and the mass is m = 10 units, find the acceleration, a. c. If the force is F = 25 units and the acceleration is a = 5 units, find the mass, m. 2. The following formula relates three quantities: Force (F), mass (m), and acceleration (a). F = ma a. Solve this equation for a. b. If the force is F = –25 units and the mass is m = 10 units, find the acceleration, a. c. If the force is F = 25 units and the acceleration is a = 5 units

OpenStudy (carson889):

a) a = F/m b) F = -25 = ma = (10)a --> a = -25/10 = -2.5 c) F = 25 = m*a = 5m --> m = 25/5 = 5

OpenStudy (anonymous):

Thank you

OpenStudy (7thgrader):

@carson889 How you do that?

OpenStudy (7thgrader):

@lilia222 You know this?